Advertisements
Advertisements
प्रश्न
A problem in Mathematics is given to three students whose chances of solving it are `1/3, 1/4` and `1/5`. What is the probability that exactly one of them will solve it?
Advertisements
उत्तर
`"P"(bar"A"_1) = 1 - "P"("A"_1) = 1 - 1/3 = 2/3`
`"P"(bar"A"_1) = 1 - "P"("A"_2) = 1 - 1/4 = 3/4`
`"P"(bar"A"_1) = 1 - "P"("A"_3) = 1 - 1/5 = 4/5`
Probability of Exactly one student solving the problem
= `"P" [("A"_1 ∩ bar"A"_2 ∩ bar"A"_3) ∪ (bar"A"_1 ∩ "A"_2 ∩ bar"A"_3) ∪ (bar"A"_1 ∩ bar"A"_2 ∩ "A"_3)]`
= `"P"("A"_1 ∩ bar"A"_2 ∩ bar"A"_3) + "P"(bar"A"_1 ∩ "A"_2 ∩ bar"A"_3) + "P"(bar"A"_1 ∩ bar"A"_2 ∩ "A"_3)`
= `"P"("A"_1) "P"(bar"A"_2) "P"(bar"A"_3) + "P"(bar"A"_1) "P"("A"_2) "P"(bar"A"_3) + "P"(bar"A"_1) "P"(bar"A"_2) "P"("A"_3)`
= `1/3 xx 3/4 xx 4/5 + 2/3 xx 1/4 xx 4/5 + 2/3 xx 3/ xx 1/5`
= `(12 + 8 + 6)/60`
= `26/16`
= `13/30`
[Probability of exactly one student solving the problem = Probability [(A1 solving the problem and A2, A3 non solving the problem) or (A1, A3 non solving and A2 solving) or (A1, A2 solving and A3 non solving)]
APPEARS IN
संबंधित प्रश्न
In a game, a man wins Rs 5 for getting a number greater than 4 and loses Rs 1 otherwise, when a fair die is thrown. The man decided to thrown a die thrice but to quit as and when he gets a number greater than 4. Find the expected value of the amount he wins/loses
If P(A) = 0.8, P(B) = 0.5 and P(B|A) = 0.4, find P(A ∪ B)
Determine P(E|F).
Mother, father and son line up at random for a family picture
E: son on one end, F: father in middle
A fair die is rolled. Consider events E = {1, 3, 5}, F = {2, 3} and G = {2, 3, 4, 5} Find P ((E ∪ F)|G) and P ((E ∩ G)|G)
A box has 20 pens of which 2 are defective. Calculate the probability that out of 5 pens drawn one by one with replacement, at most 2 are defective.
Three cards are drawn at random (without replacement) from a well-shuffled pack of 52 playing cards. Find the probability distribution of the number of red cards. Hence, find the mean of the distribution.
A bag contains 10 white balls and 15 black balls. Two balls are drawn in succession without replacement. What is the probability that, first is white and second is black?
A bag contains 10 white balls and 15 black balls. Two balls are drawn in succession without replacement. What is the probability that, one is white and other is black?
Three fair coins are tossed. What is the probability of getting three heads given that at least two coins show heads?
Select the correct option from the given alternatives :
Bag I contains 3 red and 4 black balls while another Bag II contains 5 red and 6 black balls. One ball is drawn at random from one of the bags and it is found to be red. The probability that it was drawn from Bag II
If P(A) = 0.5, P(B) = 0.8 and P(B/A) = 0.8, find P(A/B) and P(A ∪ B)
The probability that a car being filled with petrol will also need an oil change is 0.30; the probability that it needs a new oil filter is 0.40; and the probability that both the oil and filter need changing is 0.15. If the oil had to be changed, what is the probability that a new oil filter is needed?
One bag contains 5 white and 3 black balls. Another bag contains 4 white and 6 black balls. If one ball is drawn from each bag, find the probability that both are black
Choose the correct alternative:
A letter is taken at random from the letters of the word ‘ASSISTANT’ and another letter is taken at random from the letters of the word ‘STATISTICS’. The probability that the selected letters are the same is
If P(A) = 0.4, P(B) = 0.8 and P(B|A) = 0.6, then P(A ∪ B) is equal to ______.
Two cards are drawn out randomly from a pack of 52 cards one after the other, without replacement. The probability of first card being a king and second card not being a king is:
If P(A) = `1/2`, P(B) = 0, then `P(A/B)` is
Let A and B be two non-null events such that A ⊂ B. Then, which of the following statements is always correct?
If for two events A and B, P(A – B) = `1/5` and P(A) = `3/5`, then `P(B/A)` is equal to ______.
